Exercise 102


Suppose a ball is dropped from a height of 200m. If time after the ball was released is in seconds, and the height of the ball above the ground at time is given by

 , find the average velocity from to 


Show that the slope of the line joining the points A(1,1) and B is 


A line passes through the points (0, 5) and (9, −1). Find the equation of the line which is perpendicular to the line and passes through its midpoint.


Find the equation of the line that is parallel to the line y = −2x + 6 and passing through the point A(1, 10).


A square has vertices O(0, 0), A(a, 0), B(a, a) and C(0, a).

i Find the midpoint of the diagonals OB and CA.

ii Find the length of a diagonal of the square and the radius of the circle in which OABC is inscribed.

Iii Find the equation of the circle inscribing the square




1
Expert's answer
2022-02-22T08:47:46-0500

1. "h(t)=200-10t^2"


"h(2)=200-10(2)^2=160"

"h(3)=200-10(3)^2=110"

The average velocity from 2s to 3s is


"v_{ave}=\dfrac{110m-160m}{3s-2s}=-50m/s"

2.


"A(1, 1), B(2, 3+h)"

"slope=\dfrac{y_B-y_A}{x_B-x_A}=\dfrac{3+h-1}{2-1}=2+h"

3. Midpoint "(9\/2, 2)"


"slope_1=\dfrac{-1-5}{9-0}=-\dfrac{2}{3}"

If line is perpendicular then


"slope_2=-1/(-\dfrac{2}{3})=\dfrac{3}{2}"

"y=\dfrac{3}{2}x+b"

"2=\dfrac{3}{2}(\dfrac{9}{2})+b=>b=-\dfrac{19}{4}"

The equation of the line is


"y=\dfrac{3}{2}x-\dfrac{19}{4}"

4.


"y=-2x+b"

"10=-2(1)+b=>b=12"

The equation of the line is


"y=-2x+10"

5.

i.


"M(a\/2, a\/2)"

ii.


"AC=BD=a\sqrt{2}, >0"

"r=a\/2, a>0"

iii.


"R=OB/2=\dfrac{\sqrt{2}a}{a}, a>0"